# Investigating a Model-Agnostic and Imputation-Free Approach for Irregularly-Sampled Multivariate Time-Series Modeling

### v2

@@ E.6 Analyzing the Impact of Forecast Horizon
E.7 Simpler Imputation Methods
Mask
Ratio
iT
+ LOCF
iT
+ Spline
iT
+ SAITS
MissTSM
96
60% 0.18 0.002 0.18 0.002 0.37 0.08 0.22 0.005
@@ E.7 Simpler Imputation Methods
We investigate the role of simple imputation strategies, such as Last Observation Carried Forward (LOCF), within the IMTS modeling pipeline for regular MTS methods. As shown in Table 11, on seasonal datasets such as ETTm2, such straightforward methods can at times outperform more complex alternatives. While this might raise questions about the relative benefits of complex imputation methods such as SAITS, a more important consideration is whether trends observed on simpler datasets reliably translate to real-world scenarios. The SAITS paper (Du, Côté, and Liu 2023) addresses this by comparing SAITS and LOCF using a standard baseline (RNN classifier) on the PhysioNet dataset, a challenging real-world benchmark with over 80% missing values, and demonstrates the clear advantage of SAITS (see Table 12). Notably, our proposed model-agnostic approach exhibits highly competitive performance at shorter horizons and shows even greater promise as the forecasting horizon increases.
E.8 Additional Results on Masking strategies
Figures 14, 15 and 16 show additional results on comparison between random and periodic masking strategies for different horizon lengths = 96, 192, 336. Please refer to the main paper for horizon length = 720.
@@ E.8 Additional Results on Masking strategies
Figure 14: Comparison of different masking methods: MCAR vs. Periodic Masking for ETTh2, ETTm2 and Weather datasets. SA stands for SAITS and SP stands for Spline. Horizon Window: 96
@@
Figure 15: Comparison of different masking methods: MCAR vs. Periodic Masking for ETTh2, ETTm2 and Weather datasets. SA stands for SAITS and SP stands for Spline. Horizon Window: 192
@@
Figure 16: Comparison of different masking methods: MCAR vs. Periodic Masking for ETTh2, ETTm2 and Weather datasets. SA stands for SAITS and SP stands for Spline. Horizon Window: 336
Dataset
Horizon
Window
iTransformer
+ MissTSM
iTransformer
+ SAITS
iTransformer
+ Spline
@@
Dataset
Horizon
Window
PatchTST
+ MissTSM
PatchTST
+ SAITS
PatchTST
+ Spline
ETTh2 96 0.3170.004 0.5030.013 0.3240.013
192 0.3770.009 0.5120.011 0.3990.017
336 0.3800.011 0.4100.012 0.4310.005
720 0.5140.033 0.4110.002 0.4360.017
ETTm2 96 0.2020.005 0.3220.045 0.1690.000
192 0.2610.002 0.3590.036 0.2270.000
336 0.3130.001 0.4080.043 0.2850.001
720 0.4200.027 0.4590.035 0.3760.001
Weather 96 0.2060.014 0.1690.001 0.2700.110
192 0.2760.027 0.2120.000 0.2870.080
336 0.3090.024 0.2630.001 0.3250.065
720 0.3400.003 0.3330.001 0.3910.060
Table 14: MSE (meanstd) for PatchTST integrated with MissTSM against different imputation methods: SAITS and Spline under 60% masking.
Dataset
Horizon
Window
PatchTST
+ MissTSM
PatchTST
+ SAITS
PatchTST
+ Spline
ETTh2 96 0.3220.004 0.5480.050 0.3170.009
192 0.3820.011 0.5610.056 0.3800.005
336 0.3840.008 0.4680.059 0.3720.007
720 0.6210.025 0.4970.085 0.4190.015
ETTm2 96 0.2130.006 0.4050.079 0.1770.009
192 0.2660.003 0.4470.086 0.2360.009
336 0.3150.004 0.4940.095 0.2930.007
720 0.4320.025 0.5290.092 0.3860.009
Weather 96 0.2040.014 0.1660.013 0.2620.122
192 0.2490.030 0.2070.009 0.2790.092
336 0.3040.026 0.2570.007 0.3170.075
720 0.3580.012 0.3290.008 0.3830.071
Table 15: MSE (meanstd) for PatchTST integrated with MissTSM against different imputation methods: SAITS and Spline under 70% masking.
E.9 Model Agnostic Experiments
Tables 13, 14, 15 presents performance of MissTSM integrated with different models - iTransformer and PatchTST, under 60% and 70% masking fractions.
@@ G.1 Comparison against noise-agnostic methods
In this section, we compare the proposed MissTSM approach against robust methods like RobustTSF (Cheng et al. 2024) which performs time-series modeling with noisy data, where missing values and anomalies are considered as a type of noise. While (Cheng et al. 2024) is an impactful work, there are some key differences compared to MissTSM, such as, (a) RobustTSF treats missing or anomalous values as anomalies and filters them out during training, while MissTSM aim to learn latent representations from all observed time-feature pairs without dropping any time-points, fully leveraging available data; (b) While RobustTSF relies on assumptions like smooth trend filtering and Dirac-weighted anomaly scores with user-set thresholds, which may not hold in practice, MissTSM avoids such assumptions, depending purely on supervision from observed data; (c) RobustTSF assumes no missing values during testing and only imputes zeros if missing. MissTSM robustly handles missing values both during training and testing; (d) While RobustTSF is mainly univariate, MissTSM is explicitly designed for multivariate time-series, capturing dependencies across variates.
@@
Fraction Horizon
RobustTSF
+ Autoformer
MissTSM
@@ G.2 Adapting MissTSM to Anomaly Detection
While handling anomalies is not the primary focus of our work, this case study was carried out to explore the possibility of our framework to be adapted to handle other data quality issues such as anomalies. The experiment was carried out on the Pooled Server Metrics (PSM) (Abdulaal, Liu, and Lancewicki 2021) dataset using the following setup - we first use an off-the-shelf anomaly detection method, Anomaly Transformer (Xu et al. 2021), to identify anomalies in the input dataset, which are then treated as missing values in the training of MissTSM. We call this combined framework: Anomaly Transformer + MissTSM. We use 134K samples for training the Anomaly Transformer and use the remaining 84K samples for training and evaluating two forecasting models, namely, Anomaly Transformer + MissTSM, and RobustTSF (with AutoFormer backbone). We specifically split the 84K samples into train, test, and validation sets using a 7:2:1 ratio. Once trained, we evaluate both MissTSM and RobustTSF on the test set ignoring anomalous points in the forecasting window (using ground-truth labels of anomalies).
Horizon
RobustTSF +
Autoformer (MSE)
Anomaly Transformer +
MissTSM (MSE)
@@
(a) PatchTST integrated with MissTSM
(b) PatchTST imputed with SAITS
(c) PatchTST imputed with Spline
Figure 18: Qualitative comparison on ETTm2 dataset at 70% missing fraction and horizon window=336
@@
(a) PatchTST integrated with MissTSM
(b) PatchTST imputed with SAITS
(c) PatchTST imputed with Spline
Figure 20: Qualitative comparison on Weather dataset at 60% missing fraction and horizon window=336
@@
Figure 21: Qualitative comparison on Weather dataset at 60% missing fraction and horizon window=96
@@
Figure 22: Qualitative comparison on Weather dataset at 70% missing fraction and horizon window=96
Appendix H Visualization
Visualizations on Lake (see Figure 17) and benchmark datasets (see Figures 18, 19, 20, 21, 22)
Appendix I Full Results
@@
ETTh2 ETTm2 Weather
Avg
Rank
96 192 336 720 96 192 336 720 96 192 336 720
No Masking
MissTSM 0.255{0.255} 0.234{0.234} 0.316{0.316} 0.305{0.305} $0.183$ 0.209{0.209} 0.261{0.261} 0.311{0.311} $0.164$ $0.210$ $0.254$ $0.324$ $1.9$
SimMTM $0.295$ $0.356$ $0.375$ $0.404$ $0.172$ $0.223$ $0.282$ $0.374$ $0.163$ $0.203$ $0.255$ $0.326$ $2.9$
PatchTST $0.274$ $0.338$ 0.330 $0.378$ $0.164$ $0.220$ $0.277$ $0.367$ 0.151{0.151} 0.196{0.196} 0.249{0.249} 0.319{0.319} $1.7$
AutoFormer $0.501$ $0.516$ $0.565$ $0.462$ $0.352$ $0.337$ $0.494$ $0.474$ $0.306$ $0.434$ $0.437$ $0.414$ $5.9$
DLinear $0.288$ $0.383$ $0.447$ $0.605$ $0.168$ $0.224$ $0.299$ $0.414$ $0.175$ $0.219$ $0.265$ $0.323$ $4.1$
iTransformer $0.304$ $0.392$ $0.425$ $0.415$ $0.176$ $0.246$ $0.289$ $0.379$ $0.163$ $0.203$ $0.256$ $0.326$ $4.5$
MCAR Masking

### v3

@@ E.4 Analyzing the Impact of Forecast Horizon
Mask Ratio iT + LOCF iT + Spline iT + SAITS MissTSM
96 60% 0.18 0.002 0.18 0.002 0.37 0.08 0.22 0.005
@@
E.5 Simpler Imputation Methods
We investigate the role of simple imputation strategies, such as Last Observation Carried Forward (LOCF), within the IMTS modeling pipeline for regular MTS methods. As shown in Table 15, on seasonal datasets such as ETTm2, such straightforward methods can at times outperform more complex alternatives.
@@ E.5 Simpler Imputation Methods
While this might raise questions about the relative benefits of complex imputation methods such as SAITS, a more important consideration is whether trends observed on simpler datasets reliably translate to real-world scenarios. The SAITS paper (Du et al., 2023) addresses this by comparing SAITS and LOCF using a standard baseline (RNN classifier) on the PhysioNet dataset, a challenging real-world benchmark with over 80% missing values, and demonstrates the clear advantage of SAITS (see Table 16). Notably, our proposed model-agnostic approach exhibits highly competitive performance at shorter horizons and shows even greater promise as the forecasting horizon increases.
E.6 Hyper-parameter Tuning of Baseline Imputation Models
Table 17: Comparison of MissTSM with iTransformer, and PatchTST trained on HP Tuned SAITS model. Under MCAR and Periodic settings.
Horizon MissTSM iTransformer (SAITS) iTransformer (HP tuned SAITS) PatchTST (SAITS) PatchTST (HP tuned SAITS)
MCAR 96 0.243 0.492 0.465 0.440 0.389
192 0.259 0.545 0.513 0.484 0.434
336 0.283 0.579 0.546 0.434 0.387
720 0.329 0.540 0.494 0.436 0.378
Periodic 96 0.246 0.691 0.456 0.581 0.389
192 0.263 0.715 0.511 0.620 0.427
336 0.301 0.763 0.552 0.592 0.394
720 0.353 0.773 0.484 0.644 0.408
To strengthen the the claim on performance gain of MissTSM, we carry out rigorous hyper-parameter tuning of the SAITS imputation model on the ETTh2 dataset under both 70% periodic and 70% MCAR missing scenarios. The hyperparameters and their ranges considered were – n_layers: [1,2,3], d_ffn = [64, 128, 256] and (d_model, n_heads) = (64, 1), (128, 2), (256, 4) resulting in 27 total combinations. We report the results of iTransformer and PatchTST in the table below. We first trained SAITS on the 27 combinations. Then we trained the downstream model (iTransformer and PatchTST) on all the imputed files to obtain 27 different error metrics. The best configuration, based on test set results, is selected and reported in the following table. The results of the tuning experiment are presented in Table 17.
The best iTransformer results are obtained with the following SAITS configurations -
MCAR: n_layers=1, d_model=64, d_ffn=128, n_heads=1
Periodic: n_layers=1, d_model=128, d_ffn=64, n_heads=2
The best PatchTST results are obtained with the following SAITS configurations –
MCAR: n_layers=3, d_model=256, d_ffn=128, n_heads=4
Periodic: n_layers=3, d_model=256, d_ffn=64, n_heads=4
Tuning the SAITS model significantly improved the downstream performance (e.g., MSE dropping from 0.773 to 0.484 for T=720 under periodic masking) for both MCAR and periodic masking strategies. This comparison demonstrates that. However, while the imputation-based baselines do improve with specific tuning, the MissTSM framework still consistently demonstrate competitive performance compared to the newly-optimized baselines across all horizons and on both masking patterns.
@@ E.6 Hyper-parameter Tuning of Baseline Imputation Models
Figure 16: Comparison of different masking methods: MCAR vs. Periodic Masking for ETTh2, ETTm2 and Weather datasets. SA stands for SAITS and SP stands for Spline. Horizon Window: 96. Lower MSE is better
@@
Figure 17: Comparison of different masking methods: MCAR vs. Periodic Masking for ETTh2, ETTm2 and Weather datasets. SA stands for SAITS and SP stands for Spline. Horizon Window: 192. Lower MSE is better
@@
Figure 18: Comparison of different masking methods: MCAR vs. Periodic Masking for ETTh2, ETTm2 and Weather datasets. SA stands for SAITS and SP stands for Spline. Horizon Window: 336. Lower MSE is better
E.7 Additional Results on Masking strategies
Figures 16, 17 and 18 show additional results on comparison between random and periodic masking strategies for different horizon lengths = 96, 192, 336. Please refer to the main paper for horizon length = 720.
Dataset Horizon Window iTransformer + MissTSM iTransformer + SAITS iTransformer + Spline
ETTh2 96 0.3500.006 0.6450.019 0.3250.013
192 0.4820.027 0.6120.016 0.3400.017
336 0.4040.009 0.5260.015 0.4310.005
720 0.4900.012 0.5050.005 0.4360.017
ETTm2 96 0.2050.001 0.3610.066 0.1760.001
192 0.2680.016 0.4100.064 0.2400.001
336 0.3250.018 0.4580.065 0.2900.002
720 1.0800.188 0.5210.063 0.3850.003
Weather 96 0.1930.002 0.1940.004 0.2020.001
192 0.2400.004 0.2300.002 0.2400.002
336 0.2980.004 0.2800.001 0.2910.001
720 0.3640.003 0.3480.005 0.3570.001
Dataset Horizon Window iTransformer + MissTSM iTransformer + SAITS iTransformer + Spline
@@ E.7 Additional Results on Masking strategies
E.8 Model Agnostic Experiments
E.8.1 Transformer-based backbones
Tables 19, 19 presents performance of MissTSM integrated with iTransformer, under 60% and 70% masking fractions.
E.8.2 Non-transformer-based backbones
Tables 21 and 21 presents performance of MissTSM integrated with non-transformer backbone - LSTM, under 60% and 70% MCAR and Periodic masking fractions.
Table 20: MSE (meanstd) for LSTM integrated with MissTSM vs with different imputation methods: SAITS and Spline for 60% masking.
Dataset Horizon Window LSTM + MissTSM LSTM + SAITS LSTM + Spline
ETTh2 96 0.430.04 0.530.02 0.390.002
192 0.500.02 0.590.03 0.440.004
336 0.400.01 0.510.09 0.420.02
720 0.440.01 0.560.16 0.500.05
ETTm2 96 0.330.04 0.410.09 0.220.01
192 0.460.02 0.420.07 0.270.01
336 0.460.01 0.490.07 0.330.002
720 0.490.06 0.470.02 0.440.00
Weather 96 0.340.005 0.270.01 0.350.02
192 0.380.015 0.340.02 0.370.03
336 0.320.03 0.420.01 0.380.03
720 0.390.02 0.490.06 0.450.00
Table 21: MSE (meanstd) for LSTM integrated with MissTSM vs with different imputation methods: SAITS and Spline for 70% masking.
Dataset Horizon Window LSTM + MissTSM LSTM + SAITS LSTM + Spline
ETTh2 96 0.460.06 0.510.02 0.420.02
192 0.520.01 0.600.06 0.450.01
336 0.400.03 0.460.02 0.420.01
720 0.460.01 0.480.00 0.560.03
ETTm2 96 0.440.07 0.470.01 0.220.01
192 0.510.02 0.540.04 0.300.02
336 0.600.02 0.530.04 0.360.01
720 0.630.07 0.630.15 0.450.01
Weather 96 0.380.01 0.260.07 0.280.01
192 0.330.04 0.320.04 0.340.01
336 0.400.02 0.380.00 0.440.02
720 0.470.01 0.480.02 0.490.04
[…]
@@ G.1 Comparison against noise-agnostic methods
In this section, we compare the proposed MissTSM approach against robust methods like RobustTSF (Cheng et al., 2024) which performs time-series modeling with noisy data, where missing values and anomalies are considered as a type of noise. While Cheng et al. (2024) is an impactful work, there are some key differences compared to MissTSM, such as, (a) RobustTSF treats missing or anomalous values as anomalies and filters them out during training, while MissTSM aim to learn latent representations from all observed time-feature pairs without dropping any time-points, fully leveraging available data; (b) While RobustTSF relies on assumptions like smooth trend filtering and Dirac-weighted anomaly scores with user-set thresholds, which may not hold in practice, MissTSM avoids such assumptions, depending purely on supervision from observed data; (c) RobustTSF assumes no missing values during testing and only imputes zeros if missing. MissTSM robustly handles missing values both during training and testing; (d) While RobustTSF is mainly univariate, MissTSM is explicitly designed for multivariate time-series, capturing dependencies across variates.
@@
Fraction Horizon RobustTSF + Autoformer MissTSM
@@ G.2 Adapting MissTSM to Anomaly Detection
Horizon RobustTSF + Autoformer (MSE) Anomaly Transformer + MissTSM (MSE)
@@
While handling anomalies is not the primary focus of our work, this case study was carried out to explore the possibility of our framework to be adapted to handle other data quality issues such as anomalies. The experiment was carried out on the Pooled Server Metrics (PSM) (Abdulaal et al., 2021) dataset using the following setup - we first use an off-the-shelf anomaly detection method, Anomaly Transformer (Xu et al., 2021), to identify anomalies in the input dataset, which are then treated as missing values in the training of MissTSM. We call this combined framework: Anomaly Transformer + MissTSM. We use 134K samples for training the Anomaly Transformer and use the remaining 84K samples for training and evaluating two forecasting models, namely, Anomaly Transformer + MissTSM, and RobustTSF (with AutoFormer backbone). We specifically split the 84K samples into train, test, and validation sets using a 7:2:1 ratio. Once trained, we evaluate (see Table 25) both MissTSM and RobustTSF on the test set ignoring anomalous points in the forecasting window (using ground-truth labels of anomalies).
Appendix H Visualization
Visualizations on Lake (see Figure 19) and benchmark datasets (see Figures 20, 21, 22, 23, 24)
@@ Appendix H Visualization
(a) PatchTST + MissTSM
(b) PatchTST + SAITS
(c) PatchTST + Spline
Figure 20: Qualitative comparison on ETTm2 dataset at 70% missing fraction and horizon window = 336.
@@
(a) PatchTST + MissTSM
(b) PatchTST + SAITS
(c) PatchTST + Spline
Figure 22: Qualitative comparison on Weather dataset at 60% missing fraction and horizon window = 336.
@@
Figure 23: Qualitative comparison on Weather dataset at 60% missing fraction and horizon window=96.
@@
Figure 24: Qualitative comparison on Weather dataset at 70% missing fraction and horizon window=96.
Appendix I Summary of Results